Biscayne Bay Water Inc. bottles and distributes spring water. On May 14 of the current year, Biscayne Bay Water Inc. reacquired 2,200 shares of its common stock at $61 per share. On September 6, Biscayne Bay Water Inc. sold 1,700 of the reacquired shares at $64 per share. The remaining 500 shares were sold at $57 per share on November 30.

a. Journalize the transactions of May 14, September 6, and November 30. For a compound transaction, if an amount box does not require an entry, leave it blank.

May 14
Sept. 6
Nov. 30

b. What is the balance in Paid-In Capital from Sale of Treasury Stock on December 31 of the current year?
$  

c. Where will the balance in Paid-In Capital from Sale of Treasury Stock be reported on the balance sheet?

d. For what reasons might Biscayne Bay Water Inc. have purchased the treasury stock?
Answer #1
a
May 14 Treasury Stock 134200 =2200*61
      Cash 134200
Sept 6 Cash 108800 =1700*64
     Treasury Stock 103700 =1700*61
     Paid in Capital from sale of Treasury Stock 5100
Nov 30 Cash 28500 =500*57
Paid in Capital from sale of Treasury Stock 2000
     Treasury Stock 30500 =500*61
b
Paid-In Capital from Sale of Treasury Stock 3100 Credit =5100-2000
c
It will be reported in Stockholder's Equity section
d
Reasons for purchase of Treasury Stock:
For reissuance to employees as a bonus according to stock purchase agreements
To support the market price of the stock
To provide shares for resale to employees